OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@Kemist

OpenStudy (igreen):

P = 2L + 2W You wanna get 2W by itself. Subtract 2L: P - 2L = 2W Then you can divide 2 to both sides. \(\Large\frac{P - 2L}{2} = W\)